# Lattice Boltzmann methods for soft flowing matter

### v1

@@ I Introduction
Soft matter lives at the intersection between the three fundamental states of matter: gas, liquid, and solid. Much of its fascination draws from the fact that such intersection is all but a linear superposition of the three fundamental states, and supports instead genuinely new mechanical and rheological behavior, with important consequences for many applications in science and engineering Piazza (2012); Doi (2013); Fernandez-Nieves and Puertas (2016); Kleman and Lavrentovich (2003). The demise of linear superposition can be traced to the nonlinear interactions associated with the configurational degrees of freedom: for instance, foams are typically disordered assemblies of a gas (vapor) phase into a liquid water matrix and even though both phases are Newtonian, i.e. they respond to external load in direct proportion to the load intensity, once combined together, they develop a nonlinear and often non-local response. Given its paramount role in modern science and engineering, the quantitative study of soft matter has witnessed a burgeoning growth in the last decades Nagel (2017); Chen et al. (2010); Hamley (2003). This interest is further accrued by considering situations in which soft materials flow through confined geometries, since this gives rise to entirely new regimes and driven non-equilibrium steady states. In this review, we shall be concerned mostly with soft flowing systems characterized by a dispersed phase (droplets or bubbles) into a continuum matrix, say oil droplets in a water continuum, under strong geometrical confinement. Such specific systems have witnessed major development in the last decades mostly on account of progresses in experimental microfluidics, whereby one can control the spatial configurations of the droplets by simple experimental handles, typically the the ratio of oil to water mass flow and the geometrical setup of the microfluidic channel. By properly tuning such parameters, one can seamlessly move from dilute systems (droplet gas) to denser disordered states (droplet liquids) and finally even denser ordered ones, i.e. droplet solids with various topologies Raven and Marmottant (2009); Marmottant and Raven (2009) (see Fig.1).
@@
Many different experiments can be performed with these droplet-based states of matter, by changing the geometrical and physical setup. For instance, Tang and collaborators studied the intriguing non-equilibrium pattern formation phenomena that occur in dense emulsions flowing in tapered channels Y. et al. (2016) (see Fig.2a,b). These authors report an unexpected order in the flow of a concentrated emulsion in a tapered microfluidic channel, in the form of orchestrated sequences of dislocation nucleation and migration events giving rise to a highly ordered deformation mode. This suggests that nanocrystals could be made to deform in a more controlled manner than previously expected, and also hints at possible novel flow control and mixing strategies in droplet microfluidics. On a similar although distinct vein, using a microfluidic velocimetry technique, Goyon et al. Goyon and Bocquet (2010) characterized the flow of thin layers of concentrated emulsions confined in gaps of different thicknesses by surfaces of different roughness. These experiments show clear evidence of finite size effects in the flow behavior, which can be interpreted in terms of a non-local viscosity of practical relevance for applications involving thin layers, e.g. chemical or geometrical coatings. Yet another important area of experimental microfluidics concerns the realization of hierarchical emulsions (droplets within droplets). For instance, double emulsions are highly structured fluids consisting of emulsion drops that contain smaller droplets inside. Although double emulsions are potentially of commercial value, traditional fabrication by means of two emulsification steps leads to very ill-controlled structuring. Using a microcapillary device, experimentalists managed to fabricate double emulsions that contained multiple internal droplets in a core-shell geometry. By manipulating the properties of the fluid that makes up the shell, it is possible to manufacture encapsulation structures with a high degree of control and reproducibility Utada et al. (2005); Datta et al. (2014) (see Fig.2,c,d,e). These microfluidic technologies have also opened intriguing avenues in microbiology, for the detection of pathogens, antibiotic testing, cell migration, and motility Kaminski et al. (2016); Bray (2000).
@@
The latter, in particular, is crucial to a number of physiological and pathological processes, such as wound healing, embryonic development, and cancer metastasis. In many of such instances, cells are found to migrate through highly confined environments, such as dense tissues and interstices, dramatically impacting their morphology and mechanics. For instance, in Refs.Elacqua et al. (2018), the authors studied the dynamics of a cancer cell crossing a narrow constriction (whose design is inspired by physiologically-relevant conditions) and quantified shape deformations as the cell migrates through tiny pores (see Fig.2f,g). Similar dynamic behaviors have been shown in Refs.Davidson et al. (2015, 2014) using fibroblast cells moving in confined environments. In these respects, double emulsions may serve as a simplified model to study biological cells, where the innermost droplet would be a representation of a nucleus and the layer would mimic the cell cortex containing motor proteins, such as the actomyosin complex Choi et al. (2016); Mao et al. (2019).
@@
Under such conditions, computer simulations emerge as an invaluable tool for gathering information that is otherwise inaccessible through theoretical approaches or experimental methods. Microfluidic flows are typically simulated using the principles of continuum fluid dynamics, as it is established that within the bulk flow, distant from sources of heterogeneities, the continuum assumptions remain valid down to nanometric scales. Bocquet and Charlaix (2010). In the vicinity of external surfaces, such as solid boundaries, or in the interstitial region between approaching interfaces, the continuum assumption may either go under question on sheer physical grounds, or simply become computationally unwieldy on account of the high surface/volume ratios characterizing microfluidic multiphase or multicomponent flows, such as the ones discussed in this review.
Until some two decades ago, the standard and only option out of the continuum was the resort to atomistic simulation, e.g. molecular dynamics (MD). Unfortunately, notwithstanding major sustained progress in the field, MD still falls short of reaching the scales of interest for microfluidics, both in space and more so in time. As a figure of reference, even a multibillion MD simulations can only cover three spatial decades, say from 1 nm to 1 micron, which is clearly far below the size of most experimental devices. Multiscale MD-Fluid procedures have been developed in the last decades but they are still laborious in day-to-day operations. In the last three decades a third, alternative avenue, based on the intermediate level described by kinetic theory, has generated a number of very appealing and useful mesoscale computational methods, either based on suitably discretized lattice versions of Boltzmann kinetic theory Succi (2018) or coarse-grained versions of stochastic particle dynamics, such as Dissipative Particle Dynamics Groot and Warren (1997) and Stochastic Rotation methods Gompper et al. (2009). Even though each of these methods comes with its strengths and weaknesses, we believe it is fair to state that the Lattice Boltzmann stands out as the one featuring an especially high degree of physical flexibility and computational efficiency across the full spectrum of scales of motion. This is why this review is focused on recent developments of the lattice Boltzmann (LB) method specifically aimed at capturing the complexity of soft flowing matter states, such as the ones described above.
@@ II Basic physics of microfluidic soft flowing matter
We shall refer to a two-component immiscible confined microfluidic flow composed of two fluids A and B, say water and oil for the sake of concreteness. An example of the system under study is shown in Fig.3, where a collection of monodisperse droplets (generated by the breaking of the jet of the dispersed phase (A) by the flow of the continuous one (B) in the orifice) flows in the exit channel of a flow focusser.
Figure 3: Typical structure of flow focuser and emulsion confined within. The device consists of three inlet channels (of height $H\simeq 200\mu$m and width $W\simeq 100\mu$m) supplying dispersed (A) and continuous phases (B), a narrower orifice (C) (of height $h\simeq 100\mu$m) and large outlet channel ($H_{c}\simeq 500\mu$m). The droplets emerging from the outlet of the microfluidic channel are evaporated away while the continuum phase is solidified by shining UV light into the fluid, so that one is left with a solid substrate with a regular sequence of holes (scaffold). The challenge is to preserve the stability of the soft flowing crystal at increasing speed, so as to maximize the production of scaffolding material. The figure is reproduced with permission from Ref.Montessori et al. (2019a).
@@
The fourth term calls for some specific comments. It is intended to model ”near-contact” interactions, which arise whenever two interfaces come to a distance within the range of supramolecular forces, say around 10 nanometers. The physical origin of such near-contact forces may range from steric and depletion interactions to even Casimir-like interactions Derjaguin and Landau (1941); Verwey and Overbeek (1948); Lamoreaux (2004). However, in our case, we shall be agnostic regarding their physical basis and focus instead on the task of efficiently incorporating them into the LB solvers for soft flowing matter. Prior to turning to the numerics, a brief survey of basic physics is in order. It is often customary in microfluidics to define a reduced set of dimensionless groups accounting for the mesoscale physics of the system under scrutiny. The first one is the Reynolds number which measures the ratio between inertial and viscous forces and is defined as
@@
where $E_{nc}$ is a typical energy scale of near-contact interactions and $h$ is the film thickness separating the two interfaces. As a result, we obtain:
$\mathcal{N}=\frac{E_{nc}}{\sigma D^{2}}\frac{D^{4}}{h^{4}},$ (6)
where $D$ is the diameter of the droplet. By taking $D\sim 100$ $\mu$m, $h\sim 10$ nanometers and $E_{nc}\sim kT$, we obtain $N\sim\frac{0.25\cdot 10^{-20}}{70\cdot 10^{-3}\cdot 10^{-16}}\times 10^{8}\sim 10^{3}$, indicating strong dominance of near-contact forces. However, a more realistic estimate is $E_{nc}/kT\sim 10^{-3}$, leading to $N\sim O(1)$ at $h=10$ nanometers. The above relation can also be written as
@@
Due to the weak dependence on $\alpha\simeq 10^{-3}$, the above formula shows that near-contact forces prevail below $h/D\sim 10^{-4}$. For the case in point, this occurs for $h$ below 10 nanometers (since $D\sim 100\mu m$). It is also of interest to observe that such near-contact interactions require dense regimes pretty close to the maximum packing fraction $\phi_{c}$ (see Fig.4). Based on the expression
@@ II.1 The multiscale scenario and Extended Universality
In hindsight, this is attributed to a sort of Extended Universality (EU) of the physics in point, meaning by this that the rheology of the soft matter systems under investigation is to a large extent insensitive to the details of the supra-molecular scales. More precisely, the essential effects of the near-contact interactions on large-scale rheology are largely dictated by their intensity relative to capillary forces, namely the dimensionless near-contact number $\mathcal{N}$ defined in the previous section. Values of $\mathcal{N}\sim 0.1$ are sufficient to prevent coalescence and secure a regular ”crystal-like” macroscale structure Montessori et al. (2019a). Upon increasing $\mathcal{N}$ disorder sets in and the ordered crystal tends to evolve into a disordered emulsion. Of course, one cannot expect EU to hold as a general rule, but for all the applications discussed in this work, it was found to provide a pretty satisfactory description of the complex rheology of the systems under investigation. In a broad sense, one could state that the LB simulations help unveiling the EU properties of soft flowing matter where they are.
@@ III.1 Density Functional and lattice kinetic theory
The coupling between the continuity and the Navier-Stokes equations with non-ideal pressure tensor provides a self-consistent mathematical framework describing the dynamics of the binary mixture. In such a framework, the lattice kinetic theory serves as a natural bridge between the microscopic physics and the macroscopic hydrodynamic interactions. Indeed, in the kinetic theory the main ingredient is represented by the non-ideal force term $S=F[\rho]\partial_{v}f$ (we omit subscripts and vector notation for simplicity), where $f$ is the probability density function. Such a forced-streaming term can be brought to the right-hand side of the kinetic equation and treated as a soft-collision term. The possibility to handle the partial derivative in velocity space by integrating it by parts permits to move the distribution function along characteristics $\Delta{\bf x}_{v}={\bf v}\Delta t$ and include the effect of soft forces as a perturbation to the free-streaming motion of $f$.
@@
The advantage of this approach is that the streaming step (left-hand side of the equation) is exact (zero round-off error), being the distribution convected along linear characteristics, in stark contrast with the hydrodynamic formulation in which information moves along space-time dependent material lines defined by the fluid velocity. Thus, the above mathematical procedure allows the condensation of all the complex physics of moving interfaces into a purely local source term $S$. It is worth noting that, this ”perturbative” treatment works numerically as long as the magnitude of the source term is small enough to guarantee that the local Froude number is sufficiently small, i.e. $Fr=|\frac{a\Delta t}{v}|<<10^{-3}$ where $a=F/m$, a condition which may be broken in the presence of large density gradients.

### v2

# Lattice Boltzmann simulations for soft flowing matter

@@ I Introduction
Soft matter lives at the intersection between the three fundamental states of matter: gas, liquid, and solid. Much of its fascination draws from the fact that such intersection is all but a linear superposition of the three fundamental states, and supports instead genuinely new mechanical and rheological behavior, with important consequences for many applications in science and engineering Piazza (2012); Doi (2013); Fernandez-Nieves and Puertas (2016); Kleman and Lavrentovich (2003). The demise of linear superposition can be traced to the nonlinear interactions associated with the configurational degrees of freedom: for instance, foams are typically disordered assemblies of a gas (vapor) phase into a liquid water matrix and even though both phases are Newtonian, i.e. they respond to external load in direct proportion to the load intensity, once combined together, they develop a nonlinear and often non-local response. Given its paramount role in modern science and engineering, the quantitative study of soft matter has witnessed a burgeoning growth in the last decades Nagel (2017); Chen et al. (2010); Hamley (2003). This interest is further accrued by considering situations in which soft materials flow through confined geometries, since this gives rise to entirely new regimes and driven non-equilibrium steady states. In this review, we shall be concerned mostly with soft flowing systems characterized by a dispersed phase (droplets or bubbles) into a continuum matrix, say oil droplets in a water continuum, under strong geometrical confinement. Such specific systems have witnessed major development in the last decades mostly on account of progresses in experimental microfluidics, whereby one can control the spatial configurations of the droplets by simple experimental handles, typically the ratio of oil to water mass flow and the geometrical setup of the microfluidic channel. By properly tuning such parameters, one can seamlessly move from dilute systems (droplet gas) to denser disordered states (droplet liquids) and finally even denser ordered ones, i.e. droplet solids with various topologies Raven and Marmottant (2009); Marmottant and Raven (2009) (see Fig.1).
@@
Many different experiments can be performed with these droplet-based states of matter, by changing the geometrical and physical setup. For instance, Tang and collaborators studied the intriguing non-equilibrium pattern formation phenomena that occur in dense emulsions flowing in tapered channels Y. et al. (2016) (see Fig.2a,b). These authors report an unexpected order in the flow of a concentrated emulsion in a tapered microfluidic channel, in the form of orchestrated sequences of dislocation nucleation and migration events giving rise to a highly ordered deformation mode. This suggests that nanocrystals could be made to deform in a more controlled manner than previously expected, and also hints at possible novel flow control and mixing strategies in droplet microfluidics. On a similar although distinct vein, using a microfluidic velocimetry technique, Goyon et al. Goyon and Bocquet (2010) characterized the flow of thin layers of concentrated emulsions confined in gaps of different thicknesses by surfaces of different roughness. These experiments show clear evidence of finite size effects in the flow behavior, which can be interpreted in terms of a non-local viscosity of practical relevance for applications involving thin layers, e.g. chemical or geometrical coatings. Yet another important area of experimental microfluidics concerns the realization of hierarchical emulsions (droplets within droplets). For instance, double emulsions are highly structured fluids consisting of emulsion drops that contain smaller droplets inside. Although double emulsions are potentially of commercial value, traditional fabrication by means of two emulsification steps leads to very ill-controlled structuring. Using a microcapillary device, experimentalists managed to fabricate double emulsions that contained multiple internal droplets in a core-shell geometry. By manipulating the properties of the fluid that makes up the shell, it is possible to manufacture encapsulation structures with a high degree of control and reproducibility Utada et al. (2005); Datta et al. (2014) (see Fig.2c,d,e). These microfluidic technologies have also opened intriguing avenues in microbiology, for the detection of pathogens, antibiotic testing, cell migration, and motility Kaminski et al. (2016); Bray (2000).
@@
The last two, in particular, are crucial to a number of physiological and pathological processes, such as wound healing, embryonic development, and cancer metastasis. In many of such instances, cells are found to migrate through highly confined environments, such as dense tissues and interstices, dramatically impacting their morphology and mechanics. For instance, in Refs.Elacqua et al. (2018), the authors studied the dynamics of a cancer cell crossing a narrow constriction (whose design is inspired by physiologically-relevant conditions) and quantified shape deformations as the cell migrates through tiny pores (see Fig.2f,g). Similar dynamic behaviors have been observed in Refs.Davidson et al. (2015, 2014) using fibroblast cells moving in confined environments. In these respects, double emulsions may serve as a simplified model to study biological cells, where the innermost droplet would be a representation of a nucleus and the layer would mimic the cell cortex containing motor proteins, such as the actomyosin complex Choi et al. (2016); Mao et al. (2019).
@@
Under such conditions, computer simulations emerge as an invaluable tool for gathering information that is otherwise inaccessible through theoretical approaches or experimental methods. Microfluidic flows are typically simulated using the principles of continuum fluid dynamics, as it is established that within the bulk flow, distant from sources of heterogeneities, the continuum assumptions remain valid down to nanometric scales Bocquet and Charlaix (2010). In the vicinity of external surfaces, such as solid boundaries, or in the interstitial region between approaching interfaces, the continuum assumption may either go under question on sheer physical grounds, or simply become computationally unwieldy on account of the high surface/volume ratios characterizing microfluidic multiphase or multicomponent flows, such as the ones discussed in this review.
Until some two decades ago, the standard and only option out of the continuum was the resort to atomistic simulation, e.g. Molecular Dynamics (MD). Unfortunately, notwithstanding major sustained progress in the field, MD still falls short of reaching the scales of interest for microfluidics, both in space and more so in time. As a figure of reference, even a multi-billion MD simulations can only cover three spatial decades, say from 1 nanometer to 1 micrometer, which is clearly far below the size of most experimental devices. Multiscale MD-Fluid procedures have been developed in the last decades but they are still laborious in day-to-day operations. In the last three decades a third, alternative avenue, based on the intermediate level described by kinetic theory, has generated a number of very appealing and useful mesoscale computational methods, either based on suitably discretized lattice versions of Boltzmann kinetic theory Succi (2018) or coarse-grained versions of stochastic particle dynamics, such as Dissipative Particle Dynamics Groot and Warren (1997) and Stochastic Rotation methods Gompper et al. (2009). Even though each of these methods comes with its strengths and weaknesses, we believe it is fair to state that the Lattice Boltzmann (LB) stands out as the one featuring an especially high degree of physical flexibility and computational efficiency across the full spectrum of scales of motion. This is why this review is focused on recent developments of the LB method specifically aimed at capturing the complexity of soft flowing matter states, such as the ones described above.
@@ II Basic physics of microfluidic soft flowing matter
We shall refer to a confined microfluidic flow composed of two immiscible fluids A and B, say water and oil for the sake of concreteness. An example of the system under study is shown in Fig.3, where a collection of monodisperse droplets (generated by the breaking of the jet of the dispersed phase (A) by the flow of the continuous one (B) in the orifice) flows in the exit channel of a flow focuser.
Figure 3: Typical structure of flow focuser and emulsion confined within. The device consists of three inlet channels (of height $H\simeq 200\mu$m and width $W\simeq 100\mu$m) supplying dispersed (A) and continuous phases (B), a narrower orifice (C) (of height $h_{c}\simeq 100\mu$m) and large outlet channel ($H_{c}\simeq 500\mu$m). The droplets emerging from the outlet of the microfluidic channel are evaporated away while the continuum phase is solidified by shining UV light into the fluid, so that one is left with a solid substrate with a regular sequence of holes (scaffold). The challenge is to preserve the stability of the soft flowing crystal at increasing speed, so as to maximize the production of scaffolding material. The figure is reproduced with permission from Ref.Montessori et al. (2019a).
@@
The fourth term calls for some specific comments. It is intended to model ”near-contact” interactions, which arise whenever two interfaces come to a distance within the range of supramolecular forces, say around 10 nanometers. The physical origin of such near-contact forces may range from steric and depletion interactions to even Casimir-like interactions Derjaguin and Landau (1941); Verwey and Overbeek (1948); Lamoreaux (2004). However, in our case, we shall be agnostic regarding their physical basis and focus instead on the task of efficiently incorporating them into the LB solvers for soft flowing matter. Prior to turning to the numerics, a brief survey of basic physics is in order.
It is often customary in microfluidics to define a reduced set of dimensionless groups accounting for the mesoscale physics of the system under scrutiny. The first one is the Reynolds number which measures the ratio between inertial and viscous forces and is defined as
@@
where $D$ is the droplet diameter, $E_{nc}$ is a typical energy scale of near-contact interactions and $h$ is the film thickness separating the two interfaces. As a result, we obtain
$\mathcal{N}=\frac{E_{nc}}{\sigma D^{2}}\frac{D^{4}}{h^{4}}.$ (6)
By taking $D\sim 100$ $\mu$m, $h\sim 10$ nanometers and $E_{nc}\sim kT$, we obtain ${\cal N}\sim\frac{0.25\cdot 10^{-20}}{70\cdot 10^{-3}\cdot 10^{-16}}\times 10^{8}\sim 10^{3}$, indicating strong dominance of near-contact forces. However, a more realistic estimate is $E_{nc}/kT\sim 10^{-3}$, leading to ${\cal N}\sim O(1)$ at $h=10$ nanometers. The above relation can also be written as
@@
Due to the weak dependence on $\alpha\simeq 10^{-3}$ Symeonidis and Karniadakis (2006); Wenbing and An-Chang (2010), the above formula shows that near-contact forces prevail below $h/D\sim 10^{-4}$. For the case in point, this occurs for $h$ below 10 nanometers (since $D\sim 100\mu m$). It is also of interest to observe that such near-contact interactions require dense regimes pretty close to the maximum packing fraction $\phi_{c}$ (see Fig.4). Based on the expression
@@ II.1 The multiscale scenario and Extended Universality
In hindsight, this is attributed to a sort of Extended Universality (EU) of the physics in point, meaning by this that the rheology of the soft matter systems under investigation is to a large extent insensitive to the details of the supra-molecular scales. This can be formally evaluated by looking, for example, at the Cahn number $Cn=w/D$ (where $w$ is the width of droplet interface) which controls the interface dynamics. For a droplet with diameter $D\simeq 100$ $\mu$m and a typical width $w\simeq 1$ nm, the physical Cahn number is $Cn\simeq 10^{-5}$. Since in LB simulations the interface is of the order of a few lattice spacings, hence a few microns, the LB Cahn number is around $Cn_{LB}\sim 10^{-2}$, much larger than its physical counterpart. However, evidence in the literature (see Ref.Magaletti et al. (2013)) shows that the interface dynamics is largely insensitive to the specific value of the Cahn number, as long as this number is well below $1$. In other words, the Cahn number is a ”sloppy” parameter. Hence, even though the prime scope of the LB simulations is the global rheology, we expect that they can provide a reasonable account also of the small-scale interface dynamics.
A related question regards the validity of our NCI treatment in the near-interface region where $w/h\sim 1$. Here we observe that this ratio takes comparable values both in actual reality and in the LB simulations, since both $w$ and $h$ are upscaled by three orders of magnitude. This is where our coarse-graining procedure makes a strong call to EU, namely that by imposing values of $w/h$ and ${\cal N}$ similar to the experimental ones, we expect the effects of the NCI physics on the rheological behavior to be captured to a satisfactory degree of accuracy.
More precisely, the essential effects of the near-contact interactions on large-scale rheology are largely dictated by their intensity relative to capillary forces, namely the dimensionless near-contact number $\mathcal{N}$ defined in the previous section. Values of $\mathcal{N}\sim 0.1$ are sufficient to prevent coalescence and secure a regular ”crystal-like” macroscale structure Montessori et al. (2019a). Upon increasing $\mathcal{N}$, disorder sets in and the ordered crystal tends to evolve into a disordered emulsion. Of course, one cannot expect EU to hold as a general rule, but for all the applications discussed in this work, it was found to provide a pretty satisfactory description of the complex rheology of the systems under investigation. In a broad sense, one could state that the LB simulations help unveiling the EU properties of soft flowing matter where they are.
@@ III.1 Density functional and lattice kinetic theory
The coupling between the continuity and the Navier-Stokes equations with non-ideal pressure tensor provides a self-consistent mathematical framework describing the dynamics of the binary mixture. In such a framework, the lattice kinetic theory serves as a natural bridge between the microscopic physics and the macroscopic hydrodynamic interactions. Indeed, in the kinetic theory the main ingredient is represented by the non-ideal force term $S=F[\rho]\partial_{u}f$ (we omit subscripts and vector notation for simplicity), where $f$ is the probability density function. Such a forced-streaming term can be brought to the right-hand side of the kinetic equation and treated as a soft-collision term. The possibility to handle the partial derivative in velocity space by integrating it by parts permits to move the distribution function along characteristics $\Delta{\bf x}_{u}={\bf u}\Delta t$ and include the effect of soft forces as a perturbation to the free-streaming motion of $f$.
@@
The advantage of this approach is that the streaming step (left-hand side of the equation) is exact (zero round-off error), being the distribution convected along linear characteristics, in stark contrast with the hydrodynamic formulation in which information moves along space-time dependent material lines defined by the fluid velocity. Thus, the above mathematical procedure allows the condensation of all the complex physics of moving interfaces into a purely local source term $S$. It is worth noting that, this ”perturbative” treatment works numerically as long as the magnitude of the source term is small enough to guarantee that the local Froude number is sufficiently small, i.e. $Fr=|\frac{a\Delta t}{u}|<<10^{-3}$ where $a=F/m$, a condition which may be broken in the presence of large density gradients.